# UniTracker: Learning Universal Whole-Body Motion Tracker for Humanoid Robots

### v1

@@ III-A Experiment Setup
We evaluate UniTracker in both simulated and real-world environments. In simulation, policies are trained using IsaacGym [makoviychuk2021isaacgymhighperformance] with 8192 parallel environments under domain randomization [tobin2017domainrandomizationtransferringdeep]. The training data is derived from the AMASS dataset [mahmood2019amassarchivemotioncapture] and filtered by PHC [luo2023perpetualhumanoidcontrolrealtime]. We further validate our approach through sim-to-sim transfer by deploying the trained policies in MuJoCo, and conduct extensive ablation and comparative studies. Performance is assessed using four key metrics: Success Rate (SR), Mean Per Keypoint Position Error (MPKPE), Velocity Distance(Vel-Dist) and Acceleration Distance(Acc-Dist). SR reflects the overall viability and stability of the policy; MPKPE respectively quantify the accuracy of keypoint tracking in the world coordinate frame; Vel-Dist quantifies the difference in joint velocities between the reference motion and the executed motion of the robot; Acc-Dist quantifies the difference in joint acceleration between the reference motion and the executed motion of the robot. For real-world evaluation, we deploy UniTracker on the Unitree G1 humanoid robot, which stands 1.3 meters tall and has 29 degrees of freedom. We control 23 DoFs by locking the 6 wrist joints.
III-B Baselines
@@
What kind of motion tracker yields the best tracking performance? We include two baselines: (1) a universal policy trained from scratch without the teacher-student architecture, and (2) a teacher-student framework using DAgger in the second stage without a CVAE. Direct comparisons with prior methods are not included, as their implementations are not publicly available. Experimental results in Table I show that UniTracker consistently outperforms both prior methods and our baselines across all evaluation metrics.
III-C Ablation Studies
Which dataset is most suitable for training a general motion tracker? To assess the impact of training data quality, we compare two versions of the dataset: one filtered using PHC [luo2023perpetualhumanoidcontrolrealtime] and one unfiltered. The PHC-filtered dataset removes highly aggressive motions that are physically infeasible for the robot. As a result, policies trained on this dataset tend to produce smoother and more stable motions. In contrast, the results in Figure 4 which training with the unfiltered dataset leads to a noticeable increase in keypoint tracking error, indicating that the policy struggles to follow the reference motions accurately. Moreover, we observe significantly higher action rates and velocity changes, suggesting that the learned policy exhibits overly aggressive behavior, making it unsuitable for deployment on real hardware.
@@ III-C Ablation Studies
In Figure 3, we evaluate the generalization capability of UniTracker in MuJoCo by testing it on a challenging lateral squat motion, which demands strong balance control from the robot. Although UniTracker cannot precisely track all keypoints of the reference motion, it successfully produces a plausible lateral squat behavior while maintaining balance. In contrast, the DAgger baseline without the CVAE fails to remain stable and results in the robot falling, highlighting the importance of latent motion modeling for generalization to unseen motions.
In Table II, we evaluate the robustness of our policy by incrementally adding noise to the observations. In this table, a noise level of 0 indicates no observation noise, while higher levels correspond to increasing noise intensity. We observe that UniTracker consistently outperforms the variant without CVAE in both Success Rate (SR) and Mean Per Keypoint Position Error (MPkPE). Moreover, UniTracker exhibits a slower degradation in performance as noise increases, indicating stronger robustness under observation perturbations.
Figure 4: The comparison of different datasets on action rate and keypoint difference
@@
What hyperparameter choices lead to the best performance in whole-body motion tracking? We conduct ablation studies on three key hyperparameters in our framework: the CVAE latent dimension, the future window size of the reference motion used in the CVAE prior, and the weight of the KL loss term. For the latent dimension, we experiment with values of 32, 64, 128, and 256. We observe that a dimension of 64 yields the best performance. Larger latent sizes make the CVAE harder to train and do not lead to further improvements. Results are shown in Table I For the future window, which controls how many future frames of reference motion are included in the prior input, we test window sizes of 1, 5, 10, and 20 frames. The results indicate that using 5 future frames provides the best trade-off between responsiveness and stability. Lastly, we tune the KL loss coefficient with values of 1.0, 0.1, 0.01, and 0.001. We find that setting the weight to 0.1 leads to the best policy performance, effectively balancing latent regularization and reconstruction quality.
III-D Downstream Applications
We evaluate two downstream applications of our framework: motion generation and video-based motion estimation. For motion generation, we use the MDM [tevet2022humanmotiondiffusionmodel] model to produce SMPL-based [SMPL:2015] motion sequences conditioned on text input. These motions are then retargeted to the G1 humanoid robot for tracking. Both MuJoCo [todorov2012mujoco] simulation and real-world results demonstrate that our policy can accurately track motions generated from text prompts. For video estimation, we record a long motion sequence using a monocular camera, and convert it to SMPL format using GVHMR [shen2024gvhmr]. The resulting motion is retargeted to the G1 robot. Experiments in both MuJoCo and the real world show that our policy successfully tracks the reconstructed motions. These two applications highlight the strong generalization capability of our policy across different types of reference inputs.
IV Related Work
IV-A Whole-Body Controller for Humanoid Robots
Whole-body control is essential for enabling humanoid robots to perform a wide variety of complex tasks. Prior to the rise of reinforcement learning, researchers primarily relied on traditional optimization-based control methods for humanoid whole-body control [1642100, chignoli2021mithumanoidrobotdesign, a799acef5ae94895a519d794f3eaacab, li2022dynamicwalkingbipedalrobots, 1166523, 4543616, 973365]. These approaches typically required explicit mathematical modeling of both the robot and its environment, followed by real-time optimization to compute the robot’s next action. However, such methods often struggle to adapt to environmental variations, resulting in limited robustness. Additionally, they impose heavy computational demands during online execution.
@@ IV-A Whole-Body Controller for Humanoid Robots
TABLE II: We evaluate the robustness of the policy by incrementally adding noise to the observations. UniTracker consistently outperforms the variant without CVAE and exhibits a slower degradation in performance as the noise level increases.
@@
To overcome these limitations, reinforcement learning (RL) has emerged as a powerful alternative, offering the ability to learn adaptive, robust control policies directly from interaction with the environment without relying on explicit modeling. Current reinforcement learning–based whole-body controllers for humanoid robots can be categorized by the source of their control signals, including teleoperation [ze2025twistteleoperatedwholebodyimitation, lu2025mobiletelevisionpredictivemotionpriors, fu2024humanplushumanoidshadowingimitation, he2024learninghumantohumanoidrealtimewholebody, qiu2025humanoidpolicyhuman, he2024omnih2ouniversaldexteroushumantohumanoid], offline motion datasets [shao2025langwbclanguagedirectedhumanoidwholebody, shi2025adversariallocomotionmotionimitation, ji2025exbody2advancedexpressivehumanoid, cheng2024expressivewholebodycontrolhumanoid], video-based motion estimation [mao2024learningmassivehumanvideos, allshire2025visualimitationenablescontextual] [qiu2025humanoidpolicyhuman, he2024omnih2ouniversaldexteroushumantohumanoid], and high-level task commands [he2025hoverversatileneuralwholebody, xue2025unifiedgeneralhumanoidwholebody]. Teleoperation involves directly controlling the humanoid robot in real time using human input, often through motion capture systems or wearable sensors, allowing the robot to mimic human movements with high fidelity. Representative works in this area include Twist [ze2025twistteleoperatedwholebodyimitation] and H2O [he2024learninghumantohumanoidrealtimewholebody], both of which adopt a two-stage teacher-student framework. The primary difference lies in the design of the policy’s observation space. Offline motion datasets consist of pre-collected human or humanoid motion sequences, which are used as references for training control policies through motion imitation. Representative works are Exbody [cheng2024expressivewholebodycontrolhumanoid] and Exbody2 [ji2025exbody2advancedexpressivehumanoid], which begins by carefully curating an offline motion dataset and then decouples upperand lower-body motions as much as possible, aiming to maintain stability in the lower body while encourage diversity and expressiveness in the upper body. In addition, the most recent work, GMT [chen2025gmt], is the first to demonstrate tracking of 8,000 motions using a single unified policy. Video-based motion estimation methods leverage visual input from videos to extract human motion data, which can then be used to guide humanoid robot control policies. This approach enables learning from large-scale, diverse motion sources without requiring direct human demonstration. A representative work is VideoMimic [mao2024learningmassivehumanvideos], which develops a real-to-sim-to-real pipeline to model both the robot and its surrounding environment. Task commands refer to high-level, sparse control signals that specify desired outcomes or goals, such as walking direction or target position, rather than detailed joint-level motions, enabling efficient whole-body control through abstraction. Representative works include Hover [he2025hoverversatileneuralwholebody]and HugWBC [xue2025unifiedgeneralhumanoidwholebody]. Hover unifies multiple control modes into a single policy, enabling seamless transitions while retaining the strengths of each mode, thus providing a robust and scalable humanoid control solution. HugWBC designs a general task and behavior command space and employs techniques such as symmetrical loss and intervention training. This enables real-world humanoid robots to perform a variety of natural gaits—including walking, jumping, and hopping.
IV-B Environmental Adaption
Environment adaptation is essential for humanoid robots to maintain robust performance amid variations and uncertainties such as changes in terrain, external disturbances, or sensor noise in real-world conditions. One widely adopted approach to enhance environment adaptation is domain randomization (DR) [tobin2017domainrandomizationtransferringdeep]. By introducing randomized variations in simulation parameters such as physical properties, textures, lighting, and sensor noise during training, DR encourages the learned policy to generalize better and become more robust to real-world uncertainties without requiring explicit modeling of every possible scenario. Another approach to handling environmental adaptation is ASAP [he2025asapaligningsimulationrealworld], which introduces a two-stage framework to bridge the dynamics gap between simulation and the real world for agile humanoid control. In the first stage, a motion tracking policy is trained in simulation. In the second stage, this policy is fine-tuned using a delta action model learned from real-world roll-out data. By compensating for simulation-to-reality discrepancies through residual actions, ASAP enables robust and expressive whole-body motions on real hardware, outperforming traditional approaches such as system identification and domain randomization.
V CONCLUSION
In this work, we present UniTracker, a unified and robust whole-body motion tracking framework for humanoid robots. Unlike previous methods that either rely on strong reference condition or lack generalization capabilities, UniTracker leverages a CVAE-based policy architecture to balance motion expressiveness, policy robustness, and real-world deployability within a single, streamlined design. We conduct extensive experiments in simulation and on real hardware. Simulation results across multiple metrics show that UniTracker significantly outperforms our baselines. Real-world experiments on the Unitree G1 robot further confirm that our policy can track a wide range of motions—including walking, kicking, turning, dancing, and squatting—using a single general policy. Finally, we validate the generalization capabilities of UniTracker through two downstream applications: text-conditioned motion generation and video-based motion estimation, both of which demonstrate that our method can track external motion sources that were never seen during training. These results highlight UniTracker’s potential as a scalable and flexible solution for expressive whole-body humanoid control in the real world.
Name Function
tracking keypoints $exp(||{p_{t}}^{ref}-p_{t}||)$
tracking joint position $exp(||{q_{t}}^{ref}-q_{t}||)$
tracking joint velocity $exp(||\dot{{q_{t}}^{ref}}-\dot{q_{t}}||)$
tracking body linear velocity $exp(||\dot{{p_{t}}^{ref}}-\dot{p_{t}}||)$
action rate $-||a_{t}-a_{t-1}||$
torque $-\tau$
slippage $-||{v_{t}}^{foot}*{F_{t}}^{contact}||$
TABLE III: Definition of Reward Functions

### v2

@@ 3.1 Experiment Setup
We evaluate UniTracker in both simulated and real-world environments. In simulation, policies are trained using IsaacGym [26] with 8192 parallel environments under domain randomization [27]. The training data is derived from the AMASS dataset [22] and filtered by PHC [25]. We further validate our approach through sim-to-sim transfer by deploying the trained policies in MuJoCo, and conduct extensive ablation and comparative studies. Performance is assessed using four key metrics: Success Rate (SR), Mean Per Keypoint Position Error (MPKPE), Velocity Distance(Vel-Dist) and Acceleration Distance(Acc-Dist). SR reflects the overall viability and stability of the policy; MPKPE respectively quantify the accuracy of keypoint tracking in the world coordinate frame; Vel-Dist quantifies the difference in joint velocities between the reference motion and the executed motion of the robot; Acc-Dist quantifies the difference in joint acceleration between the reference motion and the executed motion of the robot. For real-world evaluation, we deploy UniTracker on the Unitree G1 humanoid robot, which stands 1.3 meters tall and has 29 degrees of freedom. We control 23 DoFs by locking the 6 wrist joints.
3.2 Baselines
@@
What kind of motion tracker yields the best tracking performance? We include two baselines: (1) a universal policy trained from scratch without the teacher-student architecture, and (2) a teacher-student framework using DAgger in the second stage without a CVAE. Direct comparisons with prior methods are not included, as their implementations are not publicly available. Experimental results in Table 1 show that UniTracker consistently outperforms our baselines across all evaluation metrics.
3.3 Ablation Studies
@@
Figure 3: Generalization Capability of UniTracker in MuJoCo
In Figure 3, we evaluate the generalization capability of UniTracker in MuJoCo by testing it on a challenging lateral squat motion, which demands strong balance control from the robot. Although UniTracker cannot precisely track all keypoints of the reference motion, it successfully produces a plausible lateral squat behavior while maintaining balance. In contrast, the DAgger baseline without the CVAE fails to remain stable and results in the robot falling, highlighting the importance of latent motion modeling for generalization to unseen motions. We have uploaded more real-world OOD videos on our website, showcasing the generalization ability of our policy.
In Table 2, we evaluate the robustness of our policy by incrementally adding noise to the observations. In this table, a noise level of 0 indicates no observation noise, while higher levels correspond to increasing noise intensity. We observe that UniTracker consistently outperforms the variant without CVAE in both Success Rate (SR) and Mean Per Keypoint Position Error (MPkPE). Moreover, UniTracker exhibits a slower degradation in performance as noise increases, indicating stronger robustness under observation perturbations.
In Figure 5, we evaluate the global awareness of our CVAE-based framework using a walking sequence in the MuJoCo simulator. The results show that, under our framework, the robot is able to walk in a straight line while closely following the reference trajectory. In contrast, the framework without CVAE gradually deviates from the reference path, demonstrating the superiority of our approach in maintaining global consistency.
@@
What hyperparameter choices lead to the best performance in whole-body motion tracking? We conduct ablation studies on three key hyperparameters in our framework: the CVAE latent dimension, the future window size of the reference motion used in the CVAE prior, and the weight of the KL loss term. For the latent dimension, we experiment with values of 32, 64, 128, and 256. We observe that a dimension of 64 yields the best performance. Larger latent sizes make the CVAE harder to train and do not lead to further improvements. Results are shown in Table 1 For the future window, which controls how many future frames of reference motion are included in the prior input, we test window sizes of 1, 5, 10, and 20 frames. The results indicate that using 5 future frames provides the best trade-off between responsiveness and stability. Lastly, we tune the KL loss coefficient with values of 1.0, 0.1, 0.01, and 0.001. We find that setting the weight to 0.1 leads to the best policy performance, effectively balancing latent regularization and reconstruction quality.
3.4 Downstream Applications
We evaluate two downstream applications of our framework: motion generation and video-based motion estimation. For motion generation, we use the MDM [28] model to produce SMPL-based [23] motion sequences conditioned on text input. These motions are then retargeted to the G1 humanoid robot for tracking. Both MuJoCo [29] simulation and real-world results demonstrate that our policy can accurately track motions generated from text prompts. For video estimation, we record a long motion sequence using a monocular camera, and convert it to SMPL format using GVHMR [30]. The resulting motion is retargeted to the G1 robot. Experiments in both MuJoCo and the real world show that our policy successfully tracks the reconstructed motions. These two applications highlight the strong generalization capability of our policy across different types of reference inputs.
3.5 Light Weight Fast Adaption
We compared our fast adaptation module against a training-from-scratch baseline. In the latter, the observations of our model remain consistent with those of the second-stage student policy, and the reward function is identical to that used in the fast adaptation module. For evaluation, we selected three challenging motion sequences—Round-Horse Kick, Side Kick, and Martial Art—along with the entire AMASS test dataset. As shown in Figure 6, our fast adaptation approach converges substantially faster than training from scratch in terms of both episode length and cumulative reward. These results underscore the advantages of our second-stage universal policy and highlight the effectiveness of the pretrained model in enabling rapid adaptation.
@@ 3.5 Light Weight Fast Adaption
Table 2: We evaluate the robustness of the policy by incrementally adding noise to the observations.
4 Related Work
4.1 Whole-Body Controller for Humanoid Robots
Whole-body control is essential for enabling humanoid robots to perform a wide variety of complex tasks. Prior to the rise of reinforcement learning, researchers primarily relied on traditional optimization-based control methods for humanoid whole-body control [31, 32, 33, 34, 35, 36, 37]. These approaches typically required explicit mathematical modeling of both the robot and its environment, followed by real-time optimization to compute the robot’s next action. However, such methods often struggle to adapt to environmental variations, resulting in limited robustness. Additionally, they impose heavy computational demands during online execution.
Figure 5: Global Consistency of UniTracker in the MuJoCo Simulator
@@ 4.1 Whole-Body Controller for Humanoid Robots
To overcome these limitations, reinforcement learning (RL) has emerged as a powerful alternative, offering the ability to learn adaptive, robust control policies directly from interaction with the environment without relying on explicit modeling. Current reinforcement learning–based whole-body controllers for humanoid robots can be categorized by the source of their control signals, including teleoperation [1, 2, 3, 4, 5, 6], offline motion datasets [7, 8, 10, 11], video-based motion estimation [14, 15] [5, 6], and high-level task commands [16, 17]. Teleoperation involves directly controlling the humanoid robot in real time using human input, often through motion capture systems or wearable sensors, allowing the robot to mimic human movements with high fidelity. Representative works in this area include Twist [1] and H2O [4], both of which adopt a two-stage teacher-student framework. The primary difference lies in the design of the policy’s observation space. Offline motion datasets consist of pre-collected human or humanoid motion sequences, which are used as references for training control policies through motion imitation. Representative works are Exbody [11] and Exbody2 [10], which begins by carefully curating an offline motion dataset and then decouples upperand lower-body motions as much as possible, aiming to maintain stability in the lower body while encourage diversity and expressiveness in the upper body. In addition, the most recent work, GMT [12], is the first to demonstrate tracking of 8,000 motions using a single unified policy. Video-based motion estimation methods leverage visual input from videos to extract human motion data, which can then be used to guide humanoid robot control policies. This approach enables learning from large-scale, diverse motion sources without requiring direct human demonstration. A representative work is VideoMimic [14], which develops a real-to-sim-to-real pipeline to model both the robot and its surrounding environment. Task commands refer to high-level, sparse control signals that specify desired outcomes or goals, such as walking direction or target position, rather than detailed joint-level motions, enabling efficient whole-body control through abstraction. Representative works include Hover [16]and HugWBC [17]. Hover unifies multiple control modes into a single policy, enabling seamless transitions while retaining the strengths of each mode, thus providing a robust and scalable humanoid control solution. HugWBC designs a general task and behavior command space and employs techniques such as symmetrical loss and intervention training. This enables real-world humanoid robots to perform a variety of natural gaits—including walking, jumping, and hopping.
Figure 6: Fast Adaption of Challenging Motions
Name Function Weight
@@
[…]
tracking body angular velocity $exp(||\dot{r_{t}}-\dot{{r_{t}}^{ref}||})$ 0.5
action rate $|a_{t}-a_{t-1}$$|$ -0.5
torque $\tau -$1e-6
slippage $|{v_{t}}^{foot}*{F_{t}}^{contact}$$|$ -1.0
termination 1.0 -200.0
Table 3: Definition of Reward Functions
5 Conclusion
In this work, we present UniTracker, a unified and scalable framework for whole-body motion tracking in humanoid robots. Built upon a three-stage training pipeline, our approach begins with a privileged teacher policy that enables high-fidelity motion tracking and effective data curation. We then introduce a CVAE-based student policy that achieves robust deployment under partial observations by modeling motion diversity and implicitly incorporating global context. To further extend the system’s adaptability, we propose a lightweight residual decoder for fast adaptation to highly dynamic or out-of-distribution motions. We validate UniTracker extensively in both simulation and real-world settings using a 29-DoF Unitree G1 humanoid. Our method successfully tracks over 8,100 motion sequences with a single policy, outperforming strong teacher-student baselines in terms of accuracy, generalization, and robustness. The results demonstrate the effectiveness of combining generative modeling with hierarchical policy distillation and residual adaptation for expressive and general-purpose humanoid control.